I'm having trouble graphing this problem and finding the points to plot:
graph the function f(x) = -x^2+6

Answer by richard1234(7193)

Pick some easy numbers, such as x = -3, -2, -1, 0, 1, 2, 3.

For each value of x, plug it in to the expression for f(x), for example:

x = -3 ==> f(-3) = -(-3)^2 + 6 = -9 + 6 = -3
x = -2 ==> f(-2) = -(-2)^2 + 6 = -4 + 6 = 2
etc.

(Hint: Note that f is an even function, i.e. f(a) = f(-a))